ABC, Inc. is expected to pay dividends of $19.26 each year infinitely. If the required rate on the stock is 16.22%, what is today's price of the stock?

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

As per Gordon model, share price is given by:

Share price = D1 / k -g

where, D1 is next years' dividend = $19.26, k is the required rate of return = 16.22% and g is the growth = Since dividends will be same every year, so growth rate of dividends will be zero.

Putting the values in the above equation, we get,

Share price = 19.26 / 16.22% - 0%

Share price = 19.26 / 16.22%

Share price = 118.74